OpenStudy (anonymous):

find the exact value sin(arcsec 3x)

OpenStudy (anonymous):

let \[\theta=arcsec3x\] \[\sec \theta=3x\] then draw a right traingle .. secant is r/x so we have r=3x x=1 and x^2+y^2=r^2 1+y^2=9x^2 y=sqrt(9x^2-1) so we have r=3x. x=1, y= sqrt(9x^2-1) sin is y/r therefore sin(arcsec3x)=sqrt(9x^2-1)/3x